5. c) Find the total energy of a moving object of mass 6.3 kg and momentum 11.5 x 10 kgms^-1

1
Expert's answer
2021-12-07T20:13:41-0500

Express velocity in terms of momentum:


"v=p\/m."

The kinetic energy:


"E=\frac12mv^2=\frac12m\bigg(\frac pm\bigg)^2=\frac{p^2}{2m},\\\space\\
E=\frac{11.5·10}{2·6.3}=9.13\text{ J}."